# STATE-NAV: Stability-Aware Traversability Estimation for Bipedal Navigation on Rough Terrain

### v2

@@ Abstract
Bipedal robots have advantages in maneuvering human-centered environments, but face greater failure risk compared to other stable mobile plarforms such as wheeled or quadrupedal robots. While learning-based traversability has been widely studied for these platforms, bipedal traversability has instead relied on manually designed rules with limited consideration of locomotion stability on rough terrain. In this work, we present the first learning-based traversability estimation and risk-sensitive navigation framework for bipedal robots operating in diverse, uneven environments. TravFormer, a transformer-based neural network, is trained to predict bipedal instability with uncertainty, enabling risk-aware and adaptive planning. Based on the network, we define traversability as stability-aware command velocity—the fastest command velocity that keeps instability below a user-defined limit. This velocity-based traversability is integrated into a hierarchical planner that combines traversability-informed Rapid Random Tree Star (TravRRT*) for time-efficient planning and Model Predictive Control (MPC) for safe execution. We validate our method in MuJoCo simulation, demonstrating improved navigation performance, with enhanced robustness and time efficiency across varying terrains compared to existing methods. Project page: https://state-nav.github.io/statenav/
Index Terms:
Humanoids, legged robots, traversability, navigation, model predictive control, stability.
{strip}[Uncaptioned image]
Figure 1: Overall diagram of the proposed traversability estimation and the navigation framework. A transformer-based bipedal instability estimator, TravFormer, is trained using body-to-stance-foot angle as a self-supervised signal to generate stability-aware command velocity map from the geometric representation of the environment. This map is further used by a hierarchically integrated TravRRT* global planner and MPC local planner to safely navigate over diverse rough terrain.
@@ I Introduction
With the rapid advancement of legged robotics, humanoid robots have garnered significant attention for their ability to navigate complex environments with rough terrain that are often inaccessible to wheeled or tracked mobile systems [1, 2]. Despite their advantages, humanoid robots face significant challenges in maintaining stability and balance, particularly on uneven or dynamically changing terrain [3, 4]. Unlike wheeled or quadrupedal robots, which benefit from their more statically stable ground contacts, humanoids have a higher risk of failure due to their limited support area and elevated center of mass, especially during rapid locomotion or abrupt terrain transitions. While various control methods have been proposed to address locomotion stability [5, 6, 7], traversability and navigation frameworks that account for the instability of bipedal locomotion on rough terrains remain largely underexplored.
@@
Existing works have explored bipedal traversability using manually designed rules based on geometry [8, 9, 10], providing a rough representation of terrain difficulty. However, these manually designed rules often fail to accurately reflect locomotion stability and performance. Meanwhile, for wheeled and quadrupedal robots, extensive research on traversability estimation has leveraged self-supervised learning with various locomotion features such as traction or IMU variances [11, 12, 13, 14, 15, 16, 17, 18, 19]. However, the relevance of these features to humanoid locomotion stability remains uncertain due to the fundamentally different locomotion strategies employed by bipedal robots. Furthermore, the greater risk of failure in humanoid locomotion requires them to operate under stricter constraints on motion aggressiveness, such as command velocity, to maintain stability across varying terrains. This challenge is often overlooked in traditional navigation planning but is crucial for ensuring safe and efficient humanoid navigation.
@@
To address these problems, we take a data-driven approach to evaluate various locomotion features and identify the one that best correlates with bipedal instability predicting fallover risk. A Transformer-based neural network, TravFormer, is then trained to predict the identified metric from terrain features and commanded velocity using simulated data. Instead of directly using the predicted instability as a traversability measure, we define bipedal traversability as the stability-aware command velocity, i.e., the fastest command velocity that maintains predicted instability below a user-defined threshold, and used it to construct a traversability map over the environment, as illustrated in Fig. 1.
@@
Estimating traversability as a stability-aware command velocity enables stability-prioritized navigation planning. Many traditional planning approaches take traversability as a cost term with tunable weights and jointly optimize it with other costs such as the path length  [20, 13, 8, 19, 16, 21]. However, varying path length scales across environments often necessitate re-tuning of weights. On the other hand, using stability-aware command velocity mitigates the effort of manual weight tuning, as terrain difficulty is inherently reflected in the temporal cost of traversal through a robot-specific instability threshold rather than environment-specific parameters. We leverage the advantage of the proposed traversability measure through our hierarchical Traversability-Informed Rapidly-exploring Random Tree Star (TravRRT*) [22, 23]–Model Predictive Control (MPC) navigation planning algorithm as shown in Fig. 1. The global TravRRT* planner simplifies the planning objective to navigation time optimization, while leveraging terrain traversability to guide informed sampling toward safer, more navigable regions. Additionally, the learned command is incorporated as a constraint in the Linear Inverted Pendulum Model (LIPM)-based local MPC planner, to generate an optimal motion plan that follows the global path while satisfying stability constraints.
@@
  • We conduct a comparative analysis of locomotion features to determine the best feature reflecting bipedal instability–the body-to-stance-foot angle–and use it as the most predictive instability feature of fallover risk across rough terrain.
@@
  • Stability-aware bipedal navigation across diverse environments is achieved by representing traversability as a risk-sensitive stability-aware command velocity and integrating it into a hierarchical Traversability-Informed RRT* (TravRRT*)–MPC framework.
  • We validate the proposed framework in MuJoCo simulations [24] using a bipedal robot, Digit, demonstrating that our framework enables the bipedal robot to navigate rough terrain safely over various rough terrain.
@@ II-A Self-Supervised Traversability Learning for Grounded Mobile Robots
Traversability characterizes a robot’s ability to navigate its environment by quantifying the level of difficulty for traversing a terrain [25]. In existing approaches, traversability has been estimated based on manually defined rules based on terrain geometry [21] or supervised learning with human-labeled semantics [26]. However, both manual cost designs and human-provided labels often fail to capture the complex interaction between robots and environments. More recent efforts have developed learning-based traversability estimators in a self-supervised manner, using supervisory signals derived from various proprioceptive features during locomotion, such as ground reaction wrench analysis [11], foothold scores [12], traction estimates [13, 14, 15], statistical features of sensor measurements such as IMU data [16, 17], and energy consumption metric [18, 19], or random motion success rate [19]. While these self-supervising features capture specific aspects of locomotion difficulty, they lack validation regarding their ability to represent the fallover risk in case of bipedal locomotion.
@@ II-B Traversability Estimation and Navigation Planning for Bipedal Locomotion
Traversability for bipedal locomotion on rough terrain remains largely underexplored. The works of [9, 10, 3] estimate terrain difficulty by simply calculating sagittal and lateral slopes of an area, penalizing or rejecting regions with steep inclines during navigation planning. An approach with a decent bipedal locomotion-specific consideration is developed in [8], which samples expected footstep candidates around nodes during A* planning and defines traversability as the percentage of cells with sufficiently low inclination and height proximity within the sampled area. Other works have explored learning-based traversability estimation for traversing flat, tilted surfaces based on footstep availability [27, 28], or adjusting body height to navigate low-ceiling environments [29]. However, these methods are limited to specific types of environments and do not address navigation over large-scale, more general rough terrains.
@@ II-C Navigation Planning with Traversability Estimation
Autonomous navigation systems have leveraged traversability estimation at multiple levels of the planning stack to generate safe and efficient paths. A common approach at the global planning level is to define traversability as a score or cost and combine it with other costs such as path length, enabling a trade-off between locomotion stability and navigation time [20, 13, 8, 19, 16, 21]. While incorporating traversability, the works in [20, 30, 14, 15] use Conditional Value at Risk (CVaR) to incorporate environmental uncertainty into the planning for risk-sensitive navigation.
At the local planning level, legged robots have extensively employed MPC to handle complex kinodynamic constraints and generate dynamically feasible motion plans. Some works integrate depth-based perception into the MPC framework to navigate irregular terrain [31, 32], leveraging terrain geometry during planning to guide footstep selection. Other works [20, 10] incorporate traversability costs into MPC for further refining the path plan from a global planner. While terrain information aids motion planning under fixed command velocities, these approaches often overlook the higher-level decision of selecting an appropriate velocity command for safety, risking failure in bipedal locomotion on terrain that requires cautious, moderated movement.
Figure 2: Randomly generated rough terrains in MuJoCo for training data collection and instability analysis.
@@ III-A Instability and Fallover Prediction
To identify an effective self-supervised signal to train our learning-based traversability estimator specifically for bipedal navigation, we begin by analyzing various locomotion features for their correlations to fallover risk in bipedal locomotion. We collect walking data from Digit in MuJoCo simulator on randomly generated rough terrains with varying geometries, such as slopes, bumps, and surface roughness—as shown in Fig. 2. Digit is controlled using a terrain-aware ALIP controller [10] commanded via random linear and angular velocities. The collected data is then parsed into gait cycles, and locomotion features are computed as the Root-Mean-Square (RMS) of raw signals per cycle. Each gait cycle is then labeled with a fallover event: 1 if a fall occurs within two subsequent steps, and 0 otherwise.
With this data, we evaluate the correlation between each feature and fallover events via logistic regression. Table I summarizes the results using McFadden’s $R^{2}$ [33] and AUC-ROC score [34], where $R^{2}$ around 0.4 and a score close to 1 generally indicates a strong relationship. While features commonly used in grounded mobile robots, such as IMU signals and traction, show acceptable correlations, the body-to-stance-foot angle described in Fig. 1 leads to better prediction accuracy. This result is reasonable because the bipedal robot counteracts its body perturbation by placing its foot further from the body to increase control authority [35]. Based on this finding, we adopt the body-to-stance-foot angle as the supervisory signal for training our traversability estimation network to predict bipedal instability.
@@ III-B Learning Instability
We aim to predict the body-to-stance-foot angle instability from the terrain and the robot state. In this work, we represent the geometrical features of the environment using a 2.5D elevation map [36]. For the robot state, many components such as instant joint values are highly time-varying and hence difficult to consider as decision variables in planning. However, other states, such as yaw orientation and velocity action commands, are relatively less time-varying and play a critical role when the learned traversability is later exploited in planning. Hence, we define our instability estimation model as follows:
@@
where $p_{\theta_{\delta}}(\cdot)$ is the target function to learn the relationship between $m_{\rm ego}$ and $a$, and the instability metric $\delta$. More specifically, $m_{\rm ego}$ is a 0.64 m $\times$ 0.64 m robo-centric (robot yaw-aligned) elevation map patch. $a$ is a command vector $(v,w)$ with linear velocity $v$ and angular velocity $w$.
@@
Energy (Control Effort) [18]
@@
Figure 3: Our TravFormer architecture, combines convolutional feature extraction with attention mechanisms to predict terrain traversability metrics for bipedal locomotion. The model processes a height-adjusted elevation map patch $m_{\rm ego}$ representing local terrain and a command vector $a=[v,\omega]^{T}$. During training, the model uses robot-centric map patches centered around the robot. At planning time, the global terrain map is divided into multiple local patches, which are processed in batch.
To model this function, we propose TravFormer, a neural network architecture specifically designed for traversability estimation of bipedal locomotion, as shown in Fig. 3. TravFormer leverages a ResNet-18 backbone [37] for terrain feature extraction and incorporates attention mechanisms [38] to focus on stability-critical aspects of the terrain while enabling context-aware integration of the command inputs. The Transformer Decoder treats command embeddings as queries and terrain features as keys/values to generate the final traversability predictions. In addition to the mean of instability, we also estimate the uncertainty of the prediction to implicitly reflect the uncertainty induced potentially by the unmodeled robot dynamics. However, as uncertainty serves as a dependent variable to the instability prediction, optimizing the instability and uncertainty simultaneously usually leads to unstable training [39]. Therefore, we adopt the two-phase sequential training strategy from [39] to address this training issue.
Phase 1 - Mean Squared Error (MSE) Training: In the initial phase, we train the network to minimize the MSE between predicted and ground truth values for instability. We train for 10 epochs during this phase, which we find sufficient to establish a strong foundation to ensure convergence for the subsequent uncertainty learning.
Phase 2 - Gaussian Negative Log-Likelihood (NLL) Training [40]: After initializing with weights from the MSE phase, we extend the network with an additional output head to predict both the mean $\hat{\delta}$ and standard deviation $\sigma$. The network is trained to minimize the negative log-likelihood of the ground truth values under a Gaussian distribution:
@@
where $N$ is the total number of samples in a batch and ${\delta}$ represents ground truth instability.
@@ IV Bipedal Navigation Planning
IV-A Traversability Estimation via Stability-Aware Command Velocity
While prior work often treats traversability as a tunable cost term, such trade-offs are less suitable for bipedal robots, where stability is critical. Instead, we define traversability as the fastest command velocity that satisfies a stability constraint, formulated in a risk-sensitive manner to account for uncertainty.

### v4

@@ Abstract
Bipedal robots have advantages in maneuvering human-centered environments, but face greater failure risk compared to other stable mobile platforms, such as wheeled or quadrupedal robots. While learning-based traversability has been widely studied for these platforms, bipedal traversability has instead relied on manually designed rules with limited consideration of locomotion stability on rough terrain. In this work, we present the first learning-based traversability estimation and risk-sensitive navigation framework for bipedal robots operating in diverse, uneven environments. TravFormer, a transformer-based neural network, is trained to predict bipedal instability with uncertainty, enabling risk-aware and adaptive planning. Based on the network, we define traversability as stability-aware command velocity—the fastest command velocity that keeps instability below a user-defined limit. This velocity-based traversability is integrated into a hierarchical planner that combines traversability-informed Rapid Random Tree Star (TravRRT*) for time-efficient path planning and Model Predictive Control (MPC) for safe execution. We validate our method in MuJoCo simulator and the real world, demonstrating improved stability, time efficiency, and robustness across diverse terrains compared with existing methods. Project page with code: https://state-nav.github.io/statenav/
@@ I Introduction
With the rapid advancement of legged robotics, humanoid robots have garnered significant attention for their ability to navigate complex environments with rough terrain that are often inaccessible to wheeled mobile systems [gu2025humanoid, krotkov2018darpa]. Despite their advantages, humanoid robots face significant challenges in maintaining stability and balance, particularly on uneven or dynamically changing terrain [huang2023efficient]. Unlike wheeled or quadrupedal robots, which benefit from their more statically stable ground contacts, humanoids have a higher risk of failure due to their limited support area and elevated center of mass, especially during rapid locomotion or abrupt terrain transitions. While various control methods have been proposed to address locomotion stability [kajita2003biped, gibson2022terrain], traversability and navigation frameworks that account for the instability of bipedal locomotion on rough terrains remain largely underexplored.
@@
Existing works have explored bipedal traversability using manually designed rules based on geometry [mccrory2023bipedal, muenprasitivej2024bipedal, shamsah2024terrain], providing a rough representation of terrain difficulty. However, these manually designed rules often fail to accurately reflect locomotion stability and performance. Meanwhile, for wheeled and quadrupedal robots, extensive research on traversability estimation has leveraged self-supervised learning with various locomotion features such as traction or IMU variances [wellhausen2019should, wellhausen2021rough, frey2023fast, cai2023probabilistic, cai2024evora, castro2023does, gan2022energy, yang2021real]. However, the relevance of these features to humanoid locomotion stability remains uncertain due to the fundamentally different locomotion strategies employed by bipedal robots. Furthermore, the greater risk of failure in humanoid locomotion requires them to operate under stricter constraints on motion aggressiveness, such as command velocity. This aspect is often overlooked in traditional navigation planning but is crucial for ensuring safe and efficient humanoid navigation.
@@
To address these problems, we take a data-driven approach to evaluate various locomotion features and identify the one that best correlates with bipedal instability for predicting fallover risk. A Transformer-based neural network, TravFormer, is then trained to predict the identified metric from terrain features and commanded velocity using simulated data. Instead of directly using the predicted instability as a traversability measure, we define traversability as the stability-aware command velocity, i.e., the fastest command velocity that maintains predicted instability below a user-defined threshold, which we use to construct a traversability map over the environment, as illustrated in Fig. LABEL:fig:overview.
@@
Estimating traversability as a stability-aware command velocity enables stability-prioritized navigation planning. Many traditional planning approaches take traversability as a cost term with tunable weights and jointly optimize it with other costs such as the path length [fan2021step, frey2023fast, mccrory2023bipedal, yang2021real, castro2023does, wermelinger2016navigation]. However, varying path length scales across environments often necessitate re-tuning of weights. On the other hand, using stability-aware command velocity mitigates the effort of manual weight tuning, as terrain difficulty is intrinsically encoded in the traversal time via a robot-specific instability threshold instead of environment-specific weights. We leverage this representation in a hierarchical Traversability-Informed Rapidly-exploring Random Tree Star (TravRRT*)–Model Predictive Control (MPC) navigation planning algorithm as shown in Fig. LABEL:fig:overview. The global TravRRT* planner simplifies the planning objective to navigation time optimization, while leveraging terrain traversability to guide informed sampling toward safer, more navigable regions. Additionally, the learned command is incorporated as a constraint in the Linear Inverted Pendulum Model (LIPM)-based local MPC planner, to generate an optimal motion plan that follows the global path while satisfying stability constraints.
@@
  • We conduct a comparative analysis of locomotion features to determine the best feature reflecting bipedal instability–the body-to-stance-foot angle (BSFA)–and use it as the supervisory signal guiding the neural network.
@@
  • Stability-aware bipedal navigation across diverse environments is achieved by representing traversability as a risk-sensitive stability-aware command velocity, providing robust, consistent performance without extensive gain tuning when integrated into the hierarchical TravRRT*–MPC framework.
  • We validate the proposed framework in MuJoCo simulator [todorov2012mujoco] and real-world experiments using a humanoid, Digit, demonstrating that our framework enables the robot to navigate diverse rough terrain safely and efficiently.
@@ II-A Self-Supervised Traversability Learning for Grounded Mobile Robots
Traversability characterizes a robot’s ability to navigate its environment by quantifying terrain difficulty [papadakis2013terrain]. In existing methods, traversability has been estimated based on manually defined rules based on terrain geometry [wermelinger2016navigation] or supervised learning with human-labeled semantics [guan2022ga]. However, both approaches often fail to capture the complex interaction between robots and environments. More recent efforts have developed learning-based traversability estimators in a self-supervised manner, using supervisory signals derived from various proprioceptive features during locomotion, such as ground reaction wrench analysis [wellhausen2019should], foothold scores [wellhausen2021rough], traction [frey2023fast, cai2023probabilistic, cai2024evora], statistical features of sensor measurements such as IMU data [castro2023does], and energy consumption metric [gan2022energy, yang2021real]. While these self-supervising features capture specific aspects of locomotion difficulty of wheeled and quadrupedal robots, they lack validation regarding their ability to represent the fallover risk in case of bipedal locomotion.
@@ II-B Traversability Estimation and Navigation Planning for Bipedal Locomotion
Traversability for bipedal locomotion on rough terrain remains largely underexplored. In [mccrory2023bipedal], A* planning samples candidate footsteps around nodes and defines traversability as the percentage of cells with sufficiently low inclination and height proximity within the sampled region. Other works have explored learning-based traversability estimation for traversing tilted flat surfaces based on footstep availability [lin2017humanoid, lin2021long], or adjusting body height to navigate low-ceiling environments [li2023autonomous]. These methods target specific settings and do not address large-scale navigation across general rough terrain.
@@ II-C Navigation Planning with Traversability Estimation
Autonomous navigation systems use traversability at multiple levels of the planning stack to navigate safely. A common approach at the global planning level is to define traversability as a score or cost and combine it with other costs such as path length, trading off between locomotion stability and navigation time [fan2021step, frey2023fast, mccrory2023bipedal, yang2021real, castro2023does, wermelinger2016navigation]. While incorporating traversability, the works in [fan2021step, cai2023probabilistic, cai2024evora] use Conditional Value at Risk (CVaR) to incorporate environmental uncertainty into the planning for risk-sensitive navigation.
At the local planning level, legged robots have extensively employed MPC to handle complex kinodynamic constraints and generate dynamically feasible motion plans. Depth-integrated MPC incorporates terrain geometry to guide footstep placement on uneven ground [grandia2023perceptive, acosta2023bipedal]. Other works [fan2021step, shamsah2024terrain] incorporate traversability costs into MPC for further refining the path plan from a global planner. However, these approaches typically fix command velocities and overlook high-level selection of velocity commands.
@@ III-A Instability and Fallover Prediction
To identify an effective self-supervised signal to train our learning-based traversability estimator specifically for bipedal navigation, we begin by analyzing various locomotion features for their correlations to fallover risk in bipedal locomotion. We collect walking data from Digit in MuJoCo on randomly generated rough terrains with varying geometries, such as slopes, bumps, and surface roughness. Digit is controlled using a terrain-aware ALIP controller [shamsah2024terrain], commanded with varying linear and angular velocities. The collected data is then parsed into gait cycles, and locomotion features are computed as the Root-Mean-Square (RMS) of raw signals per cycle. Each gait cycle is then labeled with a fallover event: 1 if a fall occurs within two subsequent steps, and 0 otherwise. All candidate features are normalized by their RMS during in-place stepping on flat ground prior to correlation analysis and training.
With this data, we evaluate the correlation between each feature and fallover events via logistic regression. Table I summarizes the results using McFadden’s $R^{2}$ [mcfadden2021quantitative] and AUC-ROC score [hanley1982meaning], where $R^{2}$ around 0.4 and a score close to 1 generally indicates a strong relationship. While features commonly used in grounded mobile robots, such as IMU signals and traction, show acceptable correlations, the body-to-stance-foot angle (BSFA) described in Fig. LABEL:fig:overview leads to better prediction accuracy. This result is reasonable because the bipedal robot counteracts its body perturbation by placing its foot further from the body to increase control authority [gibson2022terrain]. We refer to this per-cycle RMS as BSFA instability $\delta$. Moreover, based on our dataset, we empirically found that BSFA instability is only marginally influenced by slight variations in foot placement, indicating that the metric is robust to small contact-location errors. Based on this finding, we adopt the BSFA instability as the supervisory signal for training our traversability estimation network to predict bipedal instability.
@@ III-B Learning Instability
We aim to predict the BSFA instability from the terrain and the robot state. In this work, we represent the geometrical features of the environment using a 2.5D elevation map [miki2022elevation]. For the robot state, many components such as instant joint values are highly time-varying and hence difficult to consider as decision variables in planning. However, other states, such as velocity action commands, are relatively less time-varying and play a critical role when the learned traversability is later exploited in planning. Hence, we define our instability estimation model as follows:
@@
where $p_{\theta_{\delta}}(\cdot)$ is the target function that maps $m_{\rm ego}$ and $a$ to $\delta$. More specifically, $m_{\rm ego}$ is a 0.64 m $\times$ 0.64 m robot-centric (yaw-aligned) elevation map patch. $a$ is a command vector $(v,w)^{T}$ with linear velocity $v$ and angular velocity $w$. We model $\delta$ as a Gaussian distribution with mean $\hat{\delta}$ and standard deviation $\sigma$, where $\sigma$ represents the aleatoric uncertainty arising from perception noise and unmodeled effects of instantaneous joint and robot states, and is subsequently used for risk-sensitive planning. Note that external disturbances introduce unmodeled perturbations to the controller, analogous to sensor noise and mapping errors, and are therefore implicitly captured within the aleatoric uncertainty.
@@
Energy (Control Effort) [gan2022energy]
@@
Figure 2: TravFormer architecture combines convolutional feature extraction with attention mechanisms to predict BSFA instability. The model processes an elevation map patch $m_{\rm ego}$ representing local terrain and a command vector $a=[v,\omega]^{T}$. During training, the model uses robot-centric map patches centered on the robot. At planning time, the global terrain map is divided into multiple local patches, which are processed in batch.
To model this function, we propose TravFormer, a neural network architecture designed to predict BSFA instability from the elevation map and command, as shown in Fig. 2. TravFormer leverages a ResNet-18 backbone [renset] for terrain feature extraction and incorporates attention mechanisms [attn_is_all_u_need] to focus on stability-critical aspects of the terrain while enabling context-aware integration of the command inputs. The Transformer Decoder treats command embeddings as queries and terrain features as keys/values to generate the final traversability predictions. We use a two-phase schedule for stable joint optimization of $\hat{\delta}$ and $\sigma$, inspired from [tlio]
Phase 1 - Mean Squared Error (MSE) Training: In the initial phase, we train the network to minimize the MSE between the network prediction of BSFA instability and the labels collected in Section. III-A. We train for 10 epochs during this phase, which we find sufficient to establish a strong foundation to ensure convergence for the subsequent uncertainty learning.
Phase 2 - Gaussian Negative Log-Likelihood (NLL) Training [kendall2017uncertainties]: After initializing with weights from the MSE phase, we extend the network with an additional output head to predict both the mean $\hat{\delta}$ and standard deviation $\sigma$. The network is trained to match the predicted BSFA instability distribution with the label distribution under a Gaussian assumption:
@@
where $N$ is the total number of samples in a batch and $\delta$} denotes the target label collected in Section III-A.
@@ IV Bipedal Navigation Planning
IV-A Traversability as Stability-Aware Command Velocity
While prior work often treats traversability as a tunable cost term, such cases are less suitable for bipedal robots, where stability is critical. Instead, we define traversability as the fastest command velocity that satisfies a stability constraint, while accounting for uncertainty in a risk-sensitive manner.